f M (4, -6) is the midpoint of line RS. if S has coordinates (0, 13) find coordinates R.
what equation can i use to solve this?

Answer by cleomenius(959)

The equation you would use would be the midpoint formula,
((x1 + x2/ 2), (y1 + y2 / 2)) = midpoint.
(( 0 + x2 )/ 2 ) = 4
x2 = 8.
((13 + y2 )/ 2 ) = -6
y2 = -25.
The second point R is (8, -25).
Cleomenius.